If an independent-groups t test has seven participants in each group, then the degrees of freedom are​

a) 12

b) 7

c) 14

d) 6

If the null hypothesis is true, then the t test should be close to​

a) ±1.96

b) ±3.00

c) 0.00

d) ±1.65
Answer #1

Answer

(1) degree of freedom = n1 +n2 -2

we have n1 = n2 = 7

so, degree of freedom = 7+7-2. =12

option A

(2) null hypothesis Will always failed to reject the t statistic is close to the mean, i.e. when the t statistic is close to center value 0

therefore, option C is correct